OpenStudy (anonymous):

Prove that the series does in fact equal 1/3: 3/10+3/100+3/1000+3/10000+3/100000 + ...

OpenStudy (anonymous):

@Zarkon @myininaya can u guys help me?

OpenStudy (nikvist):

\[\frac{3}{10}+\frac{3}{100}+\frac{3}{1000}+\cdots=\frac{3/10}{1-1/10}=\frac{3/10}{9/10}=\frac{3}{9}=\frac{1}{3}\]geometric progression

OpenStudy (anonymous):

Thanks @nikvist